Solve y = 12xy – 2x, y(0) = 8 y(x) = Preview
Answer #1

y = 1204 - 2a 9(0) = 8 120 y = – 20 Inf. – S-120 dx – 64² e 642 - yectores * j - + ce Go? y (o) = 16+c=8, c = 47 = y(2) = 1 +